OpenStudy (anonymous):

I need help What is the simplest form of the number? -8 2/3 A. -9 B. -5.33 C. 19 D. 4

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

A?

OpenStudy (texaschic101):

-8 2/3 = -22/3 = -7.333

OpenStudy (anonymous):

I'm going with A